Costs per Equivalent Unit

The following information concerns production in the Baking Department for March. All direct materials are placed in process at the beginning of production.

ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Mar. 1 Bal., 9,000 units, 3/5 completed 15,570
31 Direct materials, 162,000 units 226,800 242,370
31 Direct labor 63,060 305,430
31 Factory overhead 35,472 340,902
31 Goods finished, 164,100 units 327,930 12,972
31 Bal. ? units, 4/5 completed 12,972

a. Based on the above data, determine each cost listed below. Round "cost per equivalent unit" answers to the nearest cent.

1. Direct materials cost per equivalent unit $
2. Conversion cost per equivalent unit $
3. Cost of the beginning work in process completed during March $
4. Cost of units started and completed during March $
5. Cost of the ending work in process $

b. Assuming that the direct materials cost is the same for February and March, did the conversion cost per equivalent unit increase, decrease, or remain the same in March?
Answer #1

Answer a:

1. Direct materials cost per equivalent unit 2. Conversion cost per equivalent unit 3. Cost of the beginning work in process

Working:

PRODUCTION and COST SUMMARY REPORT - FIFO METHOD Quantity Schedule and Equivalent Units Quantity Percentage Complete: Schedul

Answer b:

Cost of beginning work in progress = $15570

Direct material cost (in beginning work in progress) = 9000 * 1.40 = $12,600

Hence conversion cost per EUP in February = (15570-12600) / (9000 * 60%) = $0.55

In workings in answer a above, we get conversion cost per EUP in March =$0.60

Conversion cost per equivalent unit increases in March by (0.60 - 0.55=) $0.05 per EUP

Hence:

Conversion cost per equivalent unit increases in March